OpenStudy (anonymous):

If the planet's mass were doubled, how would its speed v1 change? 1) v1 would remain the same 2) v1 would increase by a factor of the square root of 2 3) v1 would increase by a factor of 2

OpenStudy (anonymous):

Well you can check this yourself, \[Fg = Fc \implies \frac{ Gm^2 }{ r^2 } =\frac{ mv^2 }{ r }\] What can you say?